Find range of c for which the line y=Mx+c touches parabola y^2=8(X+2)

Grade:11

1 Answers

Arun
5 years ago
Dear Aditya
 
y = m(x+2) + c -2m and if it is tangent to y2=8(x+2), then

c-2m = 2/m,

cm - 2m2 =2

2m2 - cm + 2=0

If m is real then

c2 ≥ 4.2.2

c2 ≥ 16

c ∈ (-∞, -4] U [4,∞)
 
Regards
Arun (askIITians forum expert)